PRECISE MEASUREMENT OF ^{40}CaH^{+} VIBRATIONAL TRANSITION FREQUENCY

\underline{MASATOSHI KAJITA}, Nat. Ins. Info. Comm. Tech., Koganei, Tokyo 184-8795, Japan; MI-NORI ABE, Dep. of Chemistry, Tokyo Metro. Univ., Hachioji, Tokyo 192-0397, Japan.

Small number of molecular ions in a linear trap can be sympathetically cooled with atomic ions and form a string crystal at the position, where the electric field is zero. Molecular ions in a strinc crystal are advantageous to measure the transition frequencies without Stark shift induced by the trap electric field, but it is required to localize small number of molecular ions in a single quantum state. ^{40}CaH^{+} molecular ion is advantageous to solve this problem, because (1) molecular ion with rotational constant of 141 GHz is localized in the vibrational-rotational ground state when the surrounding temperature is lower than 10 K, and (2) there is no hyperfine splitting in the J = 0 state. In this presentation, we porpose to measure the ^{40}CaH^{+} X^{1}Σ (v, N, F, M) = (0, 0, 1/2, ±1/2) → (v_{u}, 0, 1/2, ±1/2) (v_{u} = 1, 2, 3, , ,) transition with the uncertainty lower than 10^{−16}. With these transitions, Zeeman shift is less than 10^{−16}/G (given by the slight dependence of schielding effect by electron cloud on the vibrational state) and electric quadrupole shift is zero because of F = 1/2. The J = 0 → 0 transition is one-photon forbidden, and it can be observed also by Raman transition using two lasers. Stark shift induced by Raman lasers actually dominates the measurement uncertainty. When v = 0 → 1 transition is observed using Raman lasers in the 6000-15000 /cm, Stark shift with saturation power is of the order of 1.5 × 10^{−14} and it is higher for overtone transitions. With the following Raman laser frequencies, total Stark shift induced by two Raman lasers is zero.

v = 0 → 1 24527 /cm and 23079 /cm v = 0 → 2 24600 /cm and 21745 /cm v = 0 → 3 26237 /cm and 22017 /cm v = 0 → 4 25354 /cm and 19814 /cm

The ^{40}CaH^{+} X^{1}Σ (v, N, F, M) = (0, 0, 1/2, ±1/2) → (v_{u}, 0, 1/2, ±1/2) (v_{u} = 1, 2, 3, , ,) transition can be measured with the uncertainty lower than 10^{−16}, and it is useful to test the variation in the proton-to-electron mass ratio.